Suppose Home's demand curve for books is D=50-10P and its supply curve is S=10+10P.

a. Derive Home's import demand schedule

b. What would be the price of books in the absence of trade?
Answer #1

Answer

(a)

As Imports take place when domestic demand is greater than domestic supply and amount of imports = quantity demand - quantity supplied.

=> Imports = Demand - Supply

=> Imports(I) = D - S

=> Imports(I) = 50 - 10P - (10 + 10P)

=> Imports(I) = 40 - 20P ----------------------Import demand schedule

(b)

In the absence of trade price in the economy is the equilibrium price i.e. Price at which Demand = Supply

=> D = S

=> 50 - 10P = 10 + 10P

=> P = 40/20 = 2

Hence In the absence of trade, Price of books = 2